OpenStudy (anonymous):

How is the electrode potential of a galvanic cell determined?I need the equation....

OpenStudy (abb0t):

@thomaster

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

??? No one?

OpenStudy (frostbite):

You mean you need the reaction equation? For example for the Daniell-element?

OpenStudy (anonymous):

yah....

OpenStudy (frostbite):

Well I just write some stuff then:¨ Lets consider the Daniell-element (a galvanic cell) We can write it as a cell-diagram: \[\Large \sf Zn(s)lZn ^{2+}(aq)l l Cu ^{2+}(aq)l Cu(s)\] As half cells we get: \[\Large \sf Zn ^{2+}(aq) + 2e ^{-} \to Zn(s)\] \[\Large \sf Cu ^{2+}(aq)+2e _{-} \to Cu(s)\] Cell potential: Right cell minus left cell: \[\Large \sf Zn(s)+ Cu ^{2+}(aq) \to Zn ^{2+}(aq)+Cu(s)\]

OpenStudy (anonymous):

thanks.

OpenStudy (frostbite):

In general: In a galvanic cell the cathode has a higher potential than the anode: the species undergoing reduction, withdraws electrons from its electrode, so leaving a relative positive charge on it (corresponding to a high potential).

OpenStudy (frostbite):

Wait a sec: You are going to determine electrode potential. Stupid me. We use Nernst equation: \[\Large E=E ^{\Theta}-\frac{ RT }{ zF }\ln \left( \frac{ \left[ \sf red \right] }{ \left[ \sf ox \right] } \right)\]

OpenStudy (anonymous):

ow.lmao its all good:) lol thnks:)

OpenStudy (frostbite):

Where \(E^{\Theta}\) is the standard electron potential.

OpenStudy (frostbite):

electrode*